if a,b are solutions of x^2+2014x+1 and c,d are solutions of x^2-2015 x+1 find (a+c)(b+c)(a-b)(b-d)

Grade:10

1 Answers

Alfaz Khan
8 years ago
Multiple (i mean open the brackets), so after that you will get terms of ab,ac,ca....
all the non get able terms get cancelled and you will get ab,cd,(i hope a+b,c+d,..)
so use alpha+beta=-b/a and alpha*beta=c/a.
Hope u got it.
[THE TERMS MIGHT MATCH BUT THE APPROCHE IS CORRECT]